OpenStudy (anonymous):

find the sum 19x3 + (14x + 4x3) =

OpenStudy (anonymous):

what is with the 14x by its'self? Am I right to say that x is 'times by' not a variable?

OpenStudy (texaschic101):

19x^3 + 14x + 4x^3 = 23x^3 + 14x

OpenStudy (texaschic101):

oh...and you can factor that... 23x^3 + 14x = x(23x^2 + 14)